OpenStudy (anonymous):

What exponential function is the best fit for the data in the table? x f(x) 1 -4 3 -1 4 3 f(x) = 4(3)^x - 1 + 4 f(x) = 4(3)^x - 1 - 4 f(x) = one fourth(3)^x - 1 + 4 f(x) = one fourth(3)^x - 1 - 4

OpenStudy (lucaz):

this (3)x is \[\LARGE{3^x}\] ?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

sorry i always forget to put those in

OpenStudy (lucaz):

well, try to plug in 1 for every option, it doesn't work for any of them, you should get -4

OpenStudy (anonymous):

i am so lost

OpenStudy (lucaz):

the first option 4(3)^x - 1 + 4

OpenStudy (lucaz):

x=1 -> 4(3)^1-1+4 = 12+3 = 15

OpenStudy (lucaz):

2nd: 4(3)^1 - 1 - 4 = 12-5 = 7

OpenStudy (lucaz):

3rd: 1/4(3)^1 - 1 + 4 = 3/4+3=15/4

OpenStudy (lucaz):

4th: 1/4(3)^1 - 1 + 4=3/4+3=-17/4 it doesn't work for any of them =/

OpenStudy (anonymous):

well. this is a badly worded question :(

OpenStudy (lucaz):

you sure the table and the options are correct?

OpenStudy (anonymous):

i think its 3^x-1 -4

OpenStudy (lucaz):

oh, let's see..

OpenStudy (lucaz):

when x=1 it work for the 2nd and 4th option, but it doesn't work for x=3 or x=4

OpenStudy (lucaz):

well, anyway, you can solve this kind of problem plugging in the value of x and see if the function gives you f(x)

OpenStudy (lucaz):

I made a mistake! let me try it again

OpenStudy (anonymous):

haha no worries

OpenStudy (lucaz):

it really doesn't work.. x=1 makes 3^(1-1) or 3^0 wich is equal to 1, then: 4(1)+4=8 4(1)-4=0 1/4+4=17/4 1/4-4=-15/4 none of them is equal to -4 when x=1

OpenStudy (lucaz):

the way to solve this kind of problem is plugging in the value of x and see if it gives the correct f(x)

OpenStudy (anonymous):

so i would plug in 1/3or4 so itd be4(3)^1/3or4 -4 +4

OpenStudy (lucaz):

you plug in 1, 3 or 4

OpenStudy (lucaz):

the functions are \[\LARGE{ (something)3^{(x-1)}+something}\]

OpenStudy (lucaz):

?

OpenStudy (anonymous):

im thinkiong its D but im not sure

OpenStudy (lucaz):

sorry, I don't know what is going on, hope somebody can clarify this